Windswept Woodworks, Inc. Input Data (millions of dollars) 448 1,292 Accounts receivable 6,642 Cash & equivalents Common stock Cost of goods sold Depreciation expense Common stock dividends paid Interest expense Inventory Addition to retained earnings Long-term debt Notes payable Gross plant & equipment Retained earnings 1.196 746 2,486 Other current liabilities Tax rate Market price per share-year end Number of shares outstanding $ 17.50 500 million 500 million a. Interest coverage ratio (Assume that year 1 EBIT was 1.297 and year 1 interest expense was 120) Year 2 interest coverage ratio Year 1 intorest coverage ratio b. Average collection period (Assume that the accounts receivable balance was 950 on December 31 of the previous year and that year 1 saes were 2,700o days days Year 2 ACP Year 2 current ratio Year 1 current ratio
media%2Fdca%2Fdca5fa57-d715-422c-896c-8b
Answer #1

A. Interest Coverage Ratio

For a particular year,

Interest Coverage Ratio = EBIT / Interest Expenses

Year 1 EBIT = 1,297 (given), Year 1 Interest Expense = 120 (given)

Thus Year 1 ICR = 1,297/120 = 10.81

For Year 2,

Sales = 3,018, COGS = 1,500, Depreciation = Accumulated Depreciation for Year 2 - Accumulated Depreciation for Year 1 = 6,758 - 6,642 - 116

Thus EBIT = Sales - COGS - Accumulated Depreciation = 3018 - 1500 - 116 = 1,402

Interest Expense = 140

Thus Year 2 ICR = 1,402/140 = 10.01

B. Average Collection Period (ACP)

For Year 1,

Cash Received = Sales - Accounts Receivable = 2,700 - 840 = 1,860

Credit Sales = Sales - (Increase in Receivables - Decrease in Payables)

Assuming Decrease in Payables to be zero (as no information provided), we have,

Credit Sales = 1,860 - (950 - 840) = 1,750

Average accounts receivable = (950 + 840)/2 = 895

Average Receivables Turnover Ratio = Credit Sales / Average accounts receivable = 1,750/895 = 1.955

ACP = 365 / Average Receivables Turnover Ratio = 186.67 ~ 187 days

For Year 2,

Credit Sales = Cash Received - Receivables at the start of the period + Receivables at the end of the period = 236 - 840 + 1,292 = 688

Average accounts receivable = (840 + 1,292)/2 = 1,066

Average Receivables Turnover Ratio = Credit Sales / Average accounts receivable = 688/1,066 = 0.645

ACP = 365 / Average Receivables Turnover Ratio = 565.54 ~ 566 days

C. Current Ratio

For Year 1,

Current Assets = Accounts Receivable + Inventory + Cash & Equivalents = 840 + 1,036 + 138 = 2,014

Current Liabilities = Accounts Payable + Other Current Liabilities = 394 + 96 = 490

Thus, Current Ratio = Current Assets / Current Liabilities = 2,014/490 = 4.11

For Year 2,

Current Assets = Accounts Receivable + Inventory + Cash & Equivalents = 1,292 + 1,026 + 236 = 2,554

Current Liabilities = Accounts Payable + Other Current Liabilities = 448 + 116 = 564

Thus, Current Ratio = Current Assets / Current Liabilities = 2,554/564 = 4.53

D. Quick Ratio

For Year 1,

Quick Ratio = (Current Assets - Inventory) / Current Liabilities = (2,014 - 1,036)/490 = 2

For Year 2,

Quick Ratio = (Current Assets - Inventory) / Current Liabilities = (2,554 - 1,026)/564 = 2.71